Al AL no, 16 form Texas A&M University-Corpus Chri
Answer #1

answer 1)

option B)

solution:

F_g = F_e \Rightarrow gm = k_e\frac{q_1Q}{r^2}

solving for q1 :

q_1 = \frac{gmr^2}{k_eQ} = \frac{(10m/s^2)(0.003kg)(0.02m)^2}{k_e(-0.58\;nC)} = \boxed{-2.5\;\mu C}